OpenStudy (anonymous):

Says to solve the equation 3/2x + 4 = 1/3x + 8 For the answer I got x=24/7, is that right?

OpenStudy (campbell_st):

can I ask, is the question 1 \[\frac{3}{2x + 4} = \frac{1}{3x +8}\] or 2 \[\frac{3}{2} x + 4 = \frac{1}{3}x + 8\]

OpenStudy (anonymous):

2

OpenStudy (campbell_st):

ok... subtract 4 from both sides the subtract 1/3x from both sides \[(\frac{9}{6} - \frac{2}{6})x = 4\] so \[\frac{7}{6}x = 4 \] so your answer is correct.

OpenStudy (anonymous):

Thank you